# Maximum Length RLL Sequences in de Bruijn Graph

### v1

@@ 3 Construction of Sequences of Maximum Length
In this section, we will concentrate on constructing $(n,s)$-sequences of maximum length. We will show that there exist $(n,s)$-sequences which attain the upper bound of Corollary 2.12 and hence these sequences are maximum length $(n,s)$-sequences. We will design a method and an efficient algorithm to generate a large class of maximum length $(n,s)$-sequences. The algorithms which will be designed are implemented for contructing the next bit given the last $n$ constructed bits of the sequence. We start by describing the general method, continue with an algorithm to construct a very large class of maximum length $(n,s)$-sequences, but with a large amount of storage and the complexity for computing the next bit of the sequence is not linear if $n$ is large. This algorithm is quite efficient when $n$ is not large. We finish with an algorithm that requires a more reasonable storage size, but with a slightly smaller number of generated sequences, but the best possible for the amount of storage that is used. The complexity for finding the next bit of the sequence is linear in $n$. The following method and its algorithm is a modification of the method and the algorithms used frequently to generate de Bruijn sequences, such as the one in [9]. While the upper bound on the length of a maximum length $(n,s)$-sequences was based on an analysis of trails in $G_{n-1}(s)$, the $(n,s)$-sequences which will be generated are cycles in $G_{n}(s)$. The basic principle in the method is taking all the $(n,s)$-necklaces and merging them into one cycle. For a vertex $X=(x_{1},x_{2},\ldots,x_{n-1},x_{n})$ in $G_{n}$, its companion $X^{\prime}$ is defined by $X^{\prime}=(x_{1},x_{2},\ldots,x_{n-1},\bar{x}_{n})$, where $\bar{x}$ is the binary complement of $x$. Two vertex-disjoint cycles in $G_{n}$, ${\cal C}_{1}$ and ${\cal C}_{2}$, are joined together into one cycle, if they contain a pair of companion states $X$ on ${\cal C}_{1}$ and $X^{\prime}$ on ${\cal C}_{2}$. If $Y\rightarrow X$ is an edge on ${\cal C}_{1}$ and $Z\rightarrow X^{\prime}$ is an edge on ${\cal C}_{2}$, then keeping all the edges of ${\cal C}_{1}$ and ${\cal C}_{2}$ except for these two edges and adding the edges $Y\rightarrow X^{\prime}$ and $Z\rightarrow X$, will merge the two cycles into one cycle. The merging of all the cycles is done as follows. The $(n,s)$-necklaces are ordered by their weight from the one with the largest weight to the one with the smallest weight. We start with the necklace which contains only ones and continue to the necklace with a unique zero and start with step 2. At each step, we have a main cycle composed of the $(n,s)$-necklaces that were merged so far and also the remaining $(n,s)$-necklaces. At step $i$, $i\geq 2$, we choose on each $(n,s)$-necklace ${\cal X}$ with weight $n-i$ (this is the number of ones in each word of the necklace) a vertex whose representation ends with a zero, e.g., $X0$. Its companion $X1$ on a necklace ${\cal X}^{\prime}$ has weight $n-i+1$ and it is readily verified that ${\cal X}^{\prime}$ does not have a run of more than $s$ zeros, i.e., ${\cal X}^{\prime}$ in also an $(n,s)$-necklace. Hence, $X1$ is on the main cycle, so we can merge the necklace ${\cal X}$ into the main cycles and continue. This procedure ends when all the $(n,s)$-necklaces are merged into the main cycle.
To apply this procedure, we have to choose on each $(n,s)$-necklace whose weight is at most $n-2$ a vertex whose binary representation ends with a zero. For example, this vertex can be the one which is the smallest one as a binary number among all the vertices whose binary representation ends with a zero. The companion of this vertex has a weight greater by one and hence its cycle was merged before and hence each necklace can be merged in its time to the main cycle. This procedure will result in a cycle of $G_{n}$ which contains all the vertices of all the $(n,s)$-necklaces. If we want to form more than one cycle we have to choose two vertices that end in a zero, e.g., one vertex ending in a zero with the smallest binary representation and a second one ending in a zero with the largest binary value. Some necklaces have only one word ending in a zero. On such a necklace ${\cal X}$ the unique word which ends with a zero has the form $(1^{d-1}0)^{n/d}$ which implies that the length $d$ of the necklace is a divisor of $n$. On these necklaces, we cannot have two choices. Fortunately, the number of such necklaces is less than the number of divisors of $n$ and we can ignore them. We should use a binary vector whose length is the number of $(n,s)$-necklaces. This vector indicates which one of the two vertices should be taken to perform the merging of the cycles. If $K$ is the number of $(n,s)$-necklaces of weight at most $n-2$ in which there are at least two words ending in a zero, then the algorithm can be used to generate $2^{K}$ distinct maximum length $(n,s)$-sequences. There is one disadvantage in this method – $K$ is exponential in $n$ and hence the algorithm will require an exponential area of storage. To overcome this disadvantage we can choose two such vertices only for some of the $(n,s)$-necklaces, for example, all the $(n,s)$-necklaces of some given weights.
Algorithm to merge $(n,s)$-necklaces:
Choose two integers $21\leq k_{1}\leq k_{2}\leq n-$ 2. For each $k$, $k_{1}\leq k\leq k_{2}$, use an order of the $(n,s)$-necklaces (with at least two words ending in a zero) whose words have weight $k$, using any given efficient algorithm. For each such $k$ choose a binary vector $V(k)$ whose length is the number of such $(n,s)$-necklaces with weight $k$. Using the order of these necklaces, the $i$th entry of $V(k)$ (associated with the $i$th necklace in this order) is set to 0 if we merge the $i$-necklace through the word of the smallest value and is set to 1 if we merge through the word with the largest value.
Initially, $\beta_{0}=(1,1,\ldots,1)=1^{n}$, $\beta_{1}=(1,1,\ldots,1,0)=1^{n-1}0$ and $w(\beta_{1})=n-1$. Given $\beta_{i}=(b_{i},b_{i+1},\ldots,b_{i+n-1})$ and its weight $w(\beta_{i})$, proceed to produce the next bit which implies the next $n$-tuple $\beta_{i+1}=(b_{i+1},\ldots,b_{i+n-1},b_{i+n})$ as follows.
  1. (A1)
    Let $\beta^{*}_{i}=(b_{i+1},\ldots,b_{i+n-1},0)$ and $w(\beta^{*}_{i})=w(\beta_{i})-b_{i}$ be the weight of $\beta^{*}_{i}$.
  2. (A2)
    If the necklace which contains $\beta^{*}_{i}$ contains a run of more than $s$ zeros, then go to (A5). If the necklace contains exactly one word ending in a zero, then go to (A6).
  3. (A3)
    If $k_{1}\leq w(\beta^{*}_{i})\leq k_{2}$, then let $d$ be the value of the entry in $V(w(\beta^{*}_{i}))$ associated with the necklace that contains $\beta^{*}_{i}$.
  4. (A4)
    Examine the cyclic shifts of $\beta^{*}_{i}$ for the existence of a shift $X$ which end with a zero whose value is the smallest one if $d=0$ (or if $w(\beta^{*}_{i})$ is not between $k_{1}$ and $k_{2}$) or whose value is the largest one if $d=1$. If $\beta^{*}_{i}=X$, then go to (A6).
  5. (A5)
    Set $b_{i+n}:=b_{i}$, $w(\beta_{i+1}):=w(\beta_{i})$, and stop.
  6. (A6)
    Set $b_{i+n}:=\bar{b}_{i}$, $w(\beta_{i+1}):=w(\beta_{i})-b_{i}+\bar{b}_{i}$.
The correctness of the algorithm and the proof of the next theorem is done similarly to the related proof in [9].
Theorem 3.1.
  1. (a)
    For every choice of $k_{1}$ and $k_{2}$, algorithm merge $(n,s)$-necklaces produces the next bit of a maximum length $(n,s)$-sequence.
  2. (b)
    For a given choice of $k_{1}$ and $k_{2}$, there are $2^{K}$ distinct choices of vertices for merging the $(n,s)$-necklaces, where $K$ is the number of $(n,s)$-necklaces with weight between $k_{1}$ and $k_{2}$ having at least two words which end in a zero. Thus, algorithm merge$(n,s)$-necklaces can be used to produce $2^{K}$ distinct maximum length $(n,s)$-sequences.
  3. (c)
    The working space that algorithm merge $(n,s)$-necklaces requires to produce the next bit of a maximum length $(n,s)$-sequence is $3n+K$ bits and the work required to produce the next bit is at most $n$ cyclic shifts and about the same number of $n$-bits comparisons.
Other algorithms for generating de Bruijn sequences can be also adapted to merge all the $(n,s)$-necklaces for constructing $(n,s)$-sequences. The order of the necklaces can be done by some known algorithms as the one by Kociumaka, Radoszewski, and Rytter [19]. But, if $n$ is large, then the known algorithms for ordering the necklaces, including the one in [19] are not linear and it will increase the time complexity of our overall algorithm. If $n$ is small, then we can use this method as the complexity will not be an issue.
To overcome the complexity problem, assuming that $n$ is large enough, we suggest choosing in most cases the word that ends in a zero with the maximal value for merging the necklace into the main cycle, but in some cases to use other words as follows.
@@
Choose a large integer $k$ which depends on the number of maximum length $(n,s)$-sequences that we want to generate. On most $(n,s)$-necklaces the word the word ending in a zero with maximum value will be used to merge the necklace to the main cycle. On $k$ necklaces it will be another word. Each such a word is defined by $\theta(n,s,k)$ free parameters and hence $2^{k\cdot\theta(n,s,k)}$ maximum length $(n,s)$-sequences will be generated by the algorithm. The value of $\theta(n,s,k)$ will be increases as $n$ increases, or $s$ increases, or $k$ increases. Let $m=\lceil\log k\rceil$ and consider the ordered set $V=\{V(i)\}_{i=0}^{k-1}$ of $k$ distinct $(n,s)$-words in $k$ distinct $(n,s)$-necklaces, constructed as follows:
@@
    The first $(s+1)\lceil\frac{m}{s}\rceil+1$ bits of $V(i)$ contain ones at positions $(s+1)j$, $0\leq j\leq\lceil\frac{m}{s}\rceil$ (this guarantee that in these position there will not be a run of more than $s$ consecutive zeros). In the other $s\lceil\frac{m}{s}\rceil$ positions (between the positions of the ones) $V(i)$ forms the binary representation of $i$. These $(s+1)\lceil\frac{m}{s}\rceil+1$ bits are followed by a single zero.
@@
    For each $j\geq 1$, in position $\left((s+1)\lceil\frac{m}{s}\rceil+2\right)j$, where $\left((s+1)\lceil\frac{m}{s}\rceil+2\right)j\leq n-(s+1)\lceil\frac{m}{s}\rceil-3$, there is a zero. This guarantee that the requirement for the longest run of ones in (c2) is satisfied.
@@
    The rest of the positions which were not specified. These are free positions in which there are arbitrary assignments of zeros and ones.
@@
We first note that each word of $V$ is an $(n,s)$-word from an $(n,s)$-necklace since the word starts with a one (by (c1)) and ends with an isolated zero (by (c2)) and in each $s+1$ consecutive positions there is at least one one by (c1), (c2), (c3), and (c4). Note further that each $V(i)$ has a unique run of exactly $(s+1)\left(\lceil\frac{m}{s}\rceil\right)+2$ ones which ends just one position before the last position . All the other runs of ones in each $V(i)$ are shorter. These properties are guaranteed by (c2) and (c3). This implies that two such words from $V$ cannot be a shift of the other and also all the words are of full-order. In the first $(s+1)\left(\lceil\frac{m}{s}\rceil\right)+1$ bits there is the binary representation of $i$ (by (c1)) with separations of ones to satisfy the constraint of no more than $s$ consecutive zeros (again by (c1)). This will enable to find if the word that we consider is from the set $V$. There are many free parameters whose number is increased as $s$ get larger and/or if we choose larger $k$. The algorithm will have similar steps to algorithm merge $(n,s)$-necklaces. The algorithm will first try to see if the necklace of $\beta_{i}^{*}$ is a word from $V$ in which case the necklace is joined to the main cycle only if $\beta_{i}^{*}$ is the associated word of $V$. If $\beta_{i}^{*}$ is not a word in the set $V$, then it examines whether $\beta_{i}^{*}$ is word ending in a zero with the maximum value in its necklace. Note, that since the largest run of ones is at the end of the word followed by a unique zero, it follows that no $V(i)$ can be the word of maximum value end in a zero in its necklace (this longest run of ones starts the word of the maximum value in the necklace. The procedure itself is very similar, but slightly more complicated, than the one described in [9].

### v2

@@ 3 Construction of Sequences of Maximum Length
We will design a method and an efficient algorithm to generate a large class of maximum length $(n,s)$-sequences. The algorithms which will be designed are implemented for constructing the next bit (successor rule) given the last $n$ constructed bits of the sequence. Such algorithms for generating de Bruijn sequences are well documented in the literature starting with the work of Fredricksen [11, 12, 13] who was the first to consider efficient generation of a large set of these cycles. In [10] a very large set of such cycles were generated. Another efficient algorithm to generate a large set of de Bruijn sequences based on a set of primitive polynomials whose degrees are co-prime was presented by Li, Zeng, Li, Helleseth, and Li [27]. Many other algorithms with this flavor were designed later and general frameworks for algorithms to generate one sequence by different successor rules were given in [18] and [19] which extended the framework to other similar structures known as universal cycles. But, these frameworks are not designed for a construction of a large set of sequences. As there are a few strategies to generate many sequences we concentrate on one in which the number of generated sequences is $2^{K}$ when $K$ bits are stored. Each different assignment to these $K$ bits yields a different maximum length $(n,s)$-sequence. The amount of required storage is determined by the user subject to the value of $n$. We start by describing the general method and continue with an algorithm to construct a very large class of maximum length $(n,s)$-sequences. The method and its algorithm is a modification and a generalization of the one to generate de Bruijn sequences presented in [10]. While the upper bound on the length of a maximum length $(n,s)$-sequences was based on an analysis of trails in $G_{n-1}(s)$, the $(n,s)$-sequences which will be generated by this method are cycles in $G_{n}(s)$. The basic principle in the method is taking all the $(n,s)$-necklaces and merging them into one cycle. For a vertex $X=(x_{1}x_{2}~{}\cdots~{}x_{n-1}x_{n})$ in $G_{n}$, its companion $X^{\prime}$ is defined by $X^{\prime}=(x_{1}x_{2}~{}\cdots~{}x_{n-1}\bar{x}_{n})$, where $\bar{x}$ is the binary complement of $x$. Two vertex-disjoint cycles in $G_{n}$, ${\cal C}_{1}$ and ${\cal C}_{2}$, are joined together into one cycle, if they contain a pair of companion vertices $X$ on ${\cal C}_{1}$ and $X^{\prime}$ on ${\cal C}_{2}$. If $Y\rightarrow X$ is an edge on ${\cal C}_{1}$ and $Z\rightarrow X^{\prime}$ is an edge on ${\cal C}_{2}$, then keeping all the edges of ${\cal C}_{1}$ and ${\cal C}_{2}$ except for these two edges and adding the edges $Y\rightarrow X^{\prime}$ and $Z\rightarrow X$, will merge the two cycles into one cycle. This is depicted in Fig. 2.
Figure 2: Merging two cycles using companion vertices
The merging of all the cycles is done as follows. The $(n,s)$-necklaces are ordered by their weight from the one with the largest weight to the one with the smallest weight. We start with the necklace which contains only ones and continue to the necklace with a unique zero and start with step 2. At each step, we have a main cycle composed of the $(n,s)$-necklaces that were merged so far and also the remaining $(n,s)$-necklaces. At step $i$, $i\geq 2$, we choose on each $(n,s)$-necklace ${\cal X}$ with weight $n-i$ (this is the number of ones in each word of the necklace) a vertex whose representation ends with a zero, e.g., $(X0)$. Its companion $(X1)$ on a necklace ${\cal Y}$ has weight $n-i+1$ and it is readily verified that ${\cal Y}$ does not have a run of more than $s$ zeros, i.e., ${\cal Y}$ in also an $(n,s)$-necklace. Hence, $(X1)$ is on the main cycle, so we can merge the necklace ${\cal X}$ into the main cycle and continue. This procedure ends when all the $(n,s)$-necklaces are merged into the main cycle.
To apply this procedure, we have to choose on each $(n,s)$-necklace, whose weight is at most $n-2$, a vertex whose binary representation ends with a zero. For example, this vertex can be the one which is the maximum one as a binary number among all the vertices whose binary representation ends with a zero. The companion of this vertex has a weight greater by one and hence its cycle was merged before and therefore each necklace can be merged in its turn to the main cycle. This procedure will result in a cycle of $G_{n}$ which contains all the vertices of the $(n,s)$-necklaces. Algorithm 1 presents the formal steps of this method.
1
Input: Natural numbers $s,n\in\mathbb{N}$, $s<n-1$
Output: maximum length cyclic $(n,s)$-sequence
2 $i\leftarrow 0$ and $B_{i}\coloneqq b_{i}b_{i+1}~{}\cdots~{}b_{i+n-1}\leftarrow 1^{n}$
3
4while $B_{i}\neq 01~{}\cdots~{}1$ do
5       $Y\leftarrow b_{i+1}~{}\cdots~{}b_{i+n-1}0$
6       if the necklace of $Y$ contains a run of more than $s$ zeros then
7             $b_{i+n}\leftarrow b_{i}$ // implies that $b_{i+n}\leftarrow 1$
8       end if
9      else
10             $Z\leftarrow$ cyclic shift of $Y$ with the largest value
11             if $Y=Z$ then
12                   $b_{i+n}\leftarrow\bar{b}_{i}$
13             end if
14            else
15                   $b_{i+n}\leftarrow b_{i}$
16             end if
17
18       end if
19      $i\leftarrow i+1$
20 end while
21
22return $b_{0}b_{1}~{}\cdots~{}b_{i}$ (each $b_{i}$ is returned when computed)
Algorithm 1 merge the $(n,s)$-necklaces using the vertex of largest value in a necklace
If we want to form more than one cycle we have to choose on some necklaces two vertices that end in a zero. For $n$ which is not sufficiently large (for example a constant which is not too small) we will decide on these two necklaces and the two vertices using any chosen criteria. In this case, $Z$ in the algorithm will be the chosen vertex if $Y$ is on one of these chosen necklaces. Using any simple choice on a constant number of necklaces, the complexity of this algorithm is at most $O(n)$ per each bit if we use the algorithm of Booth [1] to find the cyclic shift with the largest value, which is the one used for merging in most cases (in [1] the smallest value is considered, but the smallest value in a necklace ${\cal Y}$ is the complement of the largest value in $\bar{{\cal Y}}$).
@@ 19      $i\leftarrow i+1$
If $n$ is large enough and we want to form a larger set of sequences, then we will choose the necklaces and these vertices that end with a zero based on a storage which is designed in advance and can have different choices as follows. Choose an integer $k$ which depends on the number of maximum length $(n,s)$-sequences that we want to generate. On most $(n,s)$-necklaces the word ending in a zero with maximum value will be used to merge the necklace into the main cycle. On $k$ necklaces it will be another word. Each such a word is defined by $\theta(n,s,k)$ free parameters and in total $K={k\cdot\theta(n,s,k)}$ free parameters. This will imply that $2^{K}$ maximum length $(n,s)$-sequences will be generated by the algorithm. The value of $\theta(n,s,k)$ will be increased as $n$ increases, or $s$ increases, or $k$ increases.
For example, let $m=\lceil\log k\rceil$ and consider the ordered set $V=\{V(i)\}_{i=0}^{k-1}$ of $k$ distinct $(n,s)$-words in $k$ distinct $(n,s)$-necklaces, constructed as follows:
@@
    The first $(s+1)\lceil\frac{m}{s}\rceil+1$ bits of $V(i)$ contain ones at positions $(s+1)j$, $0\leq j\leq\lceil\frac{m}{s}\rceil$ (this guarantee that in these positions there will not be a run of more than $s$ consecutive zeros). In the other $s\lceil\frac{m}{s}\rceil$ positions (between the positions of the ones) $V(i)$ forms the binary representation of $i$. These $(s+1)\lceil\frac{m}{s}\rceil+1$ bits are followed by a single zero.
@@
    For each $j\geq 1$, in position $\left((s+1)\lceil\frac{m}{s}\rceil+2\right)j-1$, where $\left((s+1)\lceil\frac{m}{s}\rceil+2\right)j\leq n-(s+1)\lceil\frac{m}{s}\rceil-3$, there is a zero. This guarantee that the requirement for the longest run of ones in (c2) is satisfied.
@@
    The rest of the positions which were not specified are free positions in which there are arbitrary assignments of zeros and ones (free parameters).
@@
We first note that each word of $V$ is an $(n,s)$-word from an $(n,s)$-necklace since the word starts with a one (by (c1)) and ends with an isolated zero (by (c2)) and in each $s+1$ consecutive positions there is at least one one by (c1), (c2), (c3), and (c4). Note further that each $V(i)$ has a unique run of exactly $(s+1)\left(\lceil\frac{m}{s}\rceil\right)+2$ ones which ends just one position before the last position. All the other runs of ones in each $V(i)$ are shorter. These properties are guaranteed by (c2) and (c3). This implies that two such words from $V$ cannot be a shift of the other and also all the words are from necklaces of full-order. In the first $(s+1)\left(\lceil\frac{m}{s}\rceil\right)+1$ bits there is the binary representation of $i$ (by (c1)) with separations of ones to satisfy the constraint of no more than $s$ consecutive zeros (again by (c1)). This will enable to find if the word that we consider is from the set $V$. There is a flexibility in the number of free parameters that can be increased as $s$ get larger and/or if we choose a larger $k$.
The algorithm will have similar steps to those of Algorithm 1. It will first try to see if $Y$ is contained in a necklace with another member of $V$. In this case the necklace is joined to the main cycle only if $Y$ is the associated word of $V$. If no word of $V$ is contained in the same necklace as $Y$, then it examines whether $Y$ is a word ending in a zero with the maximum value in its necklace. Note, that in $V$ since the largest run of ones is at the end of the word followed by a unique zero, it follows that no $V(i)$ can be the word which ends in a zero and of maximum value in its necklace (this longest run of ones starts the word of the maximum value in the necklace). Algorithm 2 presents the formal steps of the described algorithm.
1
Input: Natural numbers $s,n,k\in\mathbb{N}$, $s<n-1$ and an ordered set $V$
Output: maximum length cyclic $(n,s)$-sequence
2 $i\leftarrow 0$ and $B_{i}\coloneqq b_{i}b_{i+1}~{}\cdots~{}b_{i+n-1}\leftarrow 1^{n}$
3
4while $B_{i}\neq 01~{}\cdots~{}1$ do
5       $Y\leftarrow b_{i+1}~{}\cdots~{}b_{i+n-1}0$
6       if the necklace of $Y$ contains a run of more than $s$ zeros then
7             $b_{i+n}\leftarrow b_{i}$ // implies that $b_{i+n}\leftarrow 1$
8       end if
9      else
10             if $Y$ and some word of $V$ are on the same necklace then
11                  if Y is a vertex in $V$  then
12                        $b_{i+n}\leftarrow\bar{b}_{i}$
13                   end if
14                  else
15                        $b_{i+n}\leftarrow b_{i}$
16                   end if
17
18             end if
19            else
20                   $Z\leftarrow$ cyclic shift of $Y$ with the largest value
21                   if $Y=Z$ then
22                         $b_{i+n}\leftarrow\bar{b}_{i}$
23                   end if
24                  else
25                         $b_{i+n}\leftarrow b_{i}$
26                   end if
27
28             end if
29
30       end if
31      $i\leftarrow i+1$
32 end while
33
34return $b_{0}b_{1}~{}\cdots~{}b_{i}$ (each $b_{i}$ is returned when computed)
Algorithm 2 merge the $(n,s)$-necklaces with a set $V$
Theorem 3.1.
  1. (a)
    For a given choice of $k$, there are $2^{K}$, where $K=k\cdot\theta(n,s,k)$, distinct choices for the set $V$ of stored vertices to merge the $(n,s)$-necklaces. Thus, Algorithm 2 can be used to produce $2^{K}$ distinct maximum length $(n,s)$-sequences.
  2. (b)
    The working space that the procedure requires to produce the next bit of a maximum length $(n,s)$-sequence is $O(\textup{maximum}\{n,K\})$.
  3. (c)
    The complexity of the algorithm to find the next bit is $O(n)$.
Proof.
The proof of Theorem 3.1 is similar to the one proved in [10]. Each different choice of the ordered set $V$ implies different choices for the vertices via which the merging is performed. Different choices imply different $(n,s)$-sequences. The working space consists of the $K$ free parameters of the set $V$ and a constant number $n$ bits to store the current $n$ bits and $n$ bits to store the necklace and the current shift for comparison. Hence, the total working space is $O(\textup{maximum}\{n,K\})$.
As for the time complexity, there are three steps (lines) in the algorithm which are not trivial assignment or comparison. To compute if there is a run with more than $s$ zeros in $Y$ we scan $Y$ and sum the number of consecutive zeros along the necklace. This is done in $O(n)$ time per the $n$ bits of $Y$. To find if $Y$ and some word of $V$ are on the same necklace, first we compute the length of the largest run of ones in $Y$ in $O(n)$ time per $n$ bits. If this longest run is not of length $(s+1)\lceil\frac{m}{s}\rceil+2$, then by (c2) $Y$ and $V$ are not on the same necklace. Similarly, it there are two such runs of length $(s+1)\lceil\frac{m}{s}\rceil+2$ in $Y$, then by (c2) $Y$ and $V$ are not on the same necklace. If there is exactly one run of length $(s+1)\lceil\frac{m}{s}\rceil+2$ in $Y$, then we shift $Y$ in a way that this run with a zero before and a zero after it will be at the end of the word. Again, this is done in at most $O(n)$ time per the $n$ bits. Let $U$ be the obtained word. Now, if in the first $(s+1)\lceil\frac{m}{s}\rceil+1$ bits of $U$ there are ones as required of (c1), then the other values in these $(s+1)\lceil\frac{m}{s}\rceil+1$ bits indicated the exact entry of $V$ which might be equal $U$ (the shift of $Y$). Again, for this no more than $O(n)$ time per $n$ bits is required. We compare this entry of $V$ with $U$ to determine if $Y$ and this word of $V$ are on the same necklace. Again, this is done in $O(n)$ time per $n$ bits. Finally, to find the cyclic shift of $Y$ with the largest value in the necklace has complexity $O(n)$ due to [1]. Thus, the whole process will take no more than $O(n)$ time complexity per a computed bit. ∎
How large $n$ should be to make this algorithm effective? Recall, that $m=\lceil\log k\rceil$. The requirement of (c1) is $(s+1)\lceil\frac{m}{s}\rceil+1$ bits, (c2) requires $(s+1)\lceil\frac{m}{s}\rceil+4$ bits, (c3) requires at most $\frac{n-2(s+1)\lceil\frac{m}{s}\rceil}{(s+1)\lceil\frac{m}{s}\rceil+2}$ bits, and (c4) requires at most $\frac{n-2(s+1)\lceil\frac{m}{s}\rceil}{s+1}+1$ bits. This implies that $n$ must be very large, especially if $s$ is small or $k$ is large. For example, if $s=1$, then only as little as $\frac{n}{2}-2m-\frac{n-4m}{2m+2}-6$ bits in each word of $V$ are left for the free parameters. However, when $n$ is sufficiently large the number of generated $(n,s)$-sequences is considerably large.